# Six decades of the FitzHugh-Nagumo model: A guide through its spatio-temporal dynamics and influence across disciplines

### v3

@@ I Introduction
Figure 1: Exploring neuronal dynamics: The Hodgkin-Huxley model and its impact on subsequent research via the FitzHugh-Nagumo model. A. Depiction of the action potential across an axonal membrane, adapted from Hodgkin and Huxley’s seminal work Hodgkin and Huxley (1939), with the potential difference measured in millivolts and the external environment set as the zero potential reference. B. Illustration of ion channels within a neuronal axon, detailing the exchange of sodium and potassium ions between the extracellular medium (EM) and the intracellular medium (IM). C. Demonstration of excitable behavior as characterized by the HH model, adapted from Figure 17 from the original study Hodgkin and Huxley (1952). D. Publication and citation data gathered from the Web of Science using ”FitzHugh-Nagumo” as the search term. Note that these metrics do not include papers where the FHN model is not the primary focus, or papers published before the 1970s when the model name was first used. This shows the impact of the subsequent research using the FHN model as a simplified description of the HH model dynamics.
@@ I.1 Action potentials in neurons: the Hodgkin-Huxley model
Using voltage-clamp methods, they discovered that the permeability of the membrane for sodium and potassium was regulated independently, with the conductance depending on both time and the membrane potential. Remarkably, they provided a complete computational model for the action potential in a single cell, now known as the Hodgkin-Huxley (HH) model Hodgkin and Huxley (1952), which was a major advancement of our understanding. For their seminal work, they were awarded the Nobel Prize in Physiology or Medicine in 1963 Nob .
@@
$\begin{array}[]{ @{} l @{} }I=C_{m}\frac{dV_{m}}{dt}+\overline{g}_{K}n^{4}(V_{m}-V_{k})+\\ \\ \overline{g}_{Na}m^{3}h(V_{m}-V_{Na})+\overline{g}_{l}(V_{m}-V_{l})\\ \\[3.0pt] \displaystyle\frac{dn}{dt}=\alpha_{n}(V_{m})(1-n)-\beta_{n}(V_{m})n\\ \\[3.0pt] \displaystyle\frac{dm}{dt}=\alpha_{m}(V_{m})(1-m)-\beta_{m}(V_{m})m\\ \\[3.0pt] \displaystyle\frac{dh}{dt}=\alpha_{h}(V_{m})(1-h)-\beta_{h}(V_{m})h\end{array},$ (1)
Here, $I$ represents the current per unit area, while $\alpha_{i}$ and $\beta_{i}$ denote voltage-dependent rate constants specific to the $i$-th ion channel. $\overline{g}_{n}$ represents the maximum conductance value, and $n$, $m$, and $h$ are dimensionless probabilities ranging between 00 and $1$. These probabilities correspond to the activation states of potassium channel subunits, the activation of sodium channel subunits, and the inactivation of sodium channel subunits, respectively.
@@
The fact that the HH model correctly captures these excitable dynamics Hodgkin and Huxley (1952) was a major realization as it provided a detailed understanding of the measured action potential and the dynamical process of neural excitability (Fig. 1C). E: FitzHugh (1961); Izhikevich (2007a); Izhikevich and FitzHugh (2006); Upadhyay and Iyengar (2021); Massanés and Vicente (1999); Sherwood (2014); Siettos and Starke (2016); Deco et al. (2008) TW: Nagumo et al. (1962); Izhikevich and FitzHugh (2006); Bressloff (2014); Sherwood (2014) S: Goetze and Lai (2022); Brown et al. (2004); Stefanescu and Jirsa (2008); Chouzouris et al. (2018); Gerster et al. (2020); Chun-Biao et al. (2010); Al Beattie et al. (2024) O: FitzHugh (1961); Izhikevich (2007a); Izhikevich and FitzHugh (2006); Upadhyay and Iyengar (2021); Sherwood (2014); Longtin (1993) SP: Murray (2001) O: Goetze and Lai (2022) SB: FitzHugh (1961); Izhikevich (2007a); Upadhyay and Iyengar (2021) E: Goetze and Lai (2022) E: Clayton and Panfilov (2008); Keener and Sneyd (2009) SW: Davidenko et al. (1995); Courtemanche et al. (1990); Aliev and Panfilov (1996); Van Capelle and Durrer (1980); Pertsov et al. (1993); Nash and Panfilov (2004); Sermesant et al. (2006); Cherubini et al. (2008); Karma (1993); Alonso et al. (2003, 2016) TW: Keener (1988) O: Keener (1988) TW: Franzone et al. (2006); Keener (1988); Keener and Sneyd (2009); Biktashev (2002); Glass and Josephson (1995); Nagai et al. (2000) OP: Watanabe et al. (2004); Le and Sotiropoulos (2013); Panfilov and Keener (1995) O: Ullner et al. (2003a); Rosin et al. (2011); Romeira et al. (2016); Linares-Barranco et al. (1991); Keener (1983); Khakipoor et al. (2022); Xu et al. (2020); Liu et al. (2020); Barland et al. (2003); Li et al. (2021); Binczak et al. (2006, 2003) S: Ciszak et al. (2003); Zhang and Liao (2017); Li et al. (2021); Binczak et al. (2006) SB: Romeira et al. (2016); Xu et al. (2020); Bao et al. (2019); Zhang and Liao (2017); Liu et al. (2020); Barland et al. (2003); Binczak et al. (2006, 2003) SB: Heinrich et al. (2010); Zhang and Liao (2017); Li et al. (2021) E Rosin et al. (2011); Romeira et al. (2016); Liu et al. (2020); Barland et al. (2003); Binczak et al. (2006, 2003) O: Li et al. (2021); Takembo et al. (2019) SP: Takembo et al. (2019) OP: Takembo et al. (2019) O: Ginoux and Llibre (2015); Rotstein et al. (2012); Lindner (2004); Wiesenfeld et al. (1994); Desroches et al. (2012, 2008); Makarov et al. (2001); Pankratova et al. (2005); Baer and Erneux (1992, 1986); Kostova et al. (2004); Prokop and Gelens (2024); Prokop et al. (2024) TW: McKean Jr (1970); Liu and Van Vleck (2006); Champneys et al. (2007); Majumdar and Lahiri (2007); Yochelis et al. (2008); Zemskov et al. (2008); Zemskov and Loskutov (2008); Ei et al. (2008); Stich et al. (2009); Zemskov et al. (2015); Rinzel and Keller (1973); Yochelis et al. (2015); Ziepke et al. (2016); Tyson and Keener (1988a); Elphick et al. (1997); Krupa et al. (1997); Bär et al. (2000); Ikeda et al. (2000); Rabinovitch et al. (2001); Elphick et al. (1990); Hagberg and Meron (1994a); Tsyganov et al. (2014); Rombouts and Gelens (2020); Zaritski et al. (2004); Shepelev et al. (2017); Browne et al. (2008); Guckenheimer and Kuehn (2010, 2009); Berezovskaya et al. (2008); Zemskov et al. (2011, 2018); Zemskov and Epstein (2010); Hastings (1982); Bode (1997); Tobias and Knobloch (1998); Rombouts and Gelens (2021); Panfilov and Rudenko (1987); Saha et al. (2020); García-Ojalvo et al. (2001); Brauns and Marchetti (2024) S: Vaidyanathan (2015); Toral et al. (2003a); Ambrosio and Aziz-Alaoui (2012); Lindner (2004); Neiman et al. (1999); Schöll (2016a); Omelchenko et al. (2013); Semenova et al. (2016); Schöll (2016b); Omelchenko et al. (2015); Perc (2007, 2005a); Steur et al. (2009); Zaks et al. (2005); Schmidt et al. (2017); Gerster et al. (2020); Yanagita et al. (2005); Wang et al. (2009); Toral et al. (2003b); Yu et al. (2021); Brandstetter et al. (2010); Nikitin et al. (2019); Sawicki et al. (2019); Panchuk et al. (2013); Schülen et al. (2019); Schöll et al. (2009); Perlikowski et al. (2010); Saha and Feudel (2017) SB: Rocsoreanu et al. (2012); Lindner (2004); Baer and Erneux (1992, 1986); Kostova et al. (2004) SP: Bär et al. (2000); Ohta et al. (1989); Iqbal et al. (2017); Meron (1992); Petrich and Goldstein (1994); Goldstein et al. (1996); Vasiev (2004); Zheng and Shen (2015); Hagberg et al. (2006); Hagberg and Meron (1994b); Li et al. (2022); Kuznetsov et al. (2017); Gambino et al. (2022); Métens et al. (1997); Shoji et al. (2007) SB: Hoff et al. (2014); Brandibur and Kaslik (2022); Campbell and Waite (2001); Schülen et al. (2022); Yanagita et al. (2005); Wang et al. (2009); Burić and Todorović (2003); Zhen and Xu (2010); Achouri et al. (2022); Jia et al. (2015); Perlikowski et al. (2010); Saha and Feudel (2017) E: Lindner and Schimansky-Geier (1999); Lindner (2004) SW: Ambrosio and Aziz-Alaoui (2012); Keener (1986a); Tyson and Keener (1988a); Bär et al. (2000); Meron (1992); Panfilov and Hogeweg (1993); Winfree (1991a, b); Hagberg and Meron (1994b); Li et al. (2022); Bär and Eiswirth (1993a); Shima and Kuramoto (2004); Hakim and Karma (1999); Lin et al. (2004); Bär and Eiswirth (1993b); Dowle et al. (1997); Barkley (1991) O: Campbell and Waite (2001); Ambrosio and Aziz-Alaoui (2012); Zaks et al. (2005); Yanagita et al. (2005); Burić and Todorović (2003); Perlikowski et al. (2010); Saha and Feudel (2017); Ullner et al. (2003b) SB: Ei et al. (2008); Iqbal et al. (2017); Zheng and Shen (2015); Hagberg et al. (2006); Hagberg and Meron (1994b); Shepelev et al. (2017); Guckenheimer and Kuehn (2010, 2009); Berezovskaya et al. (2008); Tobias and Knobloch (1998); Brauns et al. (2020); Frohoff-Hülsmann and Thiele (2023) OP: Neiman et al. (1999); Schülen et al. (2022); Perc (2007); Schmidt et al. (2017); Sawicki et al. (2019) OP: Bär et al. (2000); Meron (1992); Hagberg and Meron (1994a); Vasiev (2004); Nekhamkina and Sheintuch (2006); Zaritski et al. (2004); Perc (2005b); Lin et al. (2004); Sagués et al. (2007); Brauns and Marchetti (2024) SW: Ambrosio and Aziz-Alaoui (2012); Lindner (2004); García-Ojalvo and Schimansky-Geier (1999) O: Champneys et al. (2007); Stich et al. (2009); Ambrosio and Aziz-Alaoui (2012); Rabinovitch et al. (2001); Elphick et al. (1990); Rombouts and Gelens (2020, 2021) SP: Lei and Yang (2021); Perlikowski et al. (2010) TW: Sawicki et al. (2019) O: Parra-Rivas et al. (2023a); Sneyd et al. (2017a); Hartzell et al. (2015); Tsai et al. (2008); Tyler et al. (2022) TW: Gelens et al. (2014); Noguchi (2023); Nolet et al. (2020a); Miller and Dunkel (2020); Duffy et al. (1998); Lee (1997) S: Kori et al. (2012); Cartwright (2000); Scialla et al. (2021) E: Parra-Rivas et al. (2023a); Hartzell et al. (2015) O: Gelens et al. (2014); Nolet et al. (2020a); Sneyd et al. (2017b); Lee (1997) SB: Vodeneev et al. (2011); Sukhov et al. (2011) SB: Chiangga et al. (2016); Stratford et al. (2019) SB: Gambino et al. (2023); Sneyd et al. (2017b) TW: Vodeneev et al. (2011); Sukhov et al. (2011) SP: Gambino et al. (2022, 2023) O: Kori et al. (2012) E: Noguchi (2023) SW: Lee (1997)
@@ I.2 A simplified description of neuronal excitability: the FitzHugh-Nagumo model
FitzHugh started from the van der Pol oscillator equation Kanamaru (2007), introduced in the 1920s by Van der Pol Van (1927). As its name suggests, the Van der Pol oscillator admits oscillations which are relaxation-like, meaning they feature periods of slow progress close to a ”low state” and a ”high state” punctuated by rapid transitions between them. Van der Pol’s equation is built from the simple differential equation for the damped harmonic oscillator by replacing the damping constant with a damping coefficient that depends quadratically on $x$, thus introducing a nonlinearity:
@@
$\begin{array}[]{ @{} l @{} }x_{t}=c\left[y-\left(\displaystyle\frac{x^{3}}{3}-x\right)\right],\\ \\[3.0pt] y_{t}=-\displaystyle\frac{1}{c}x,\end{array}$ (4)
@@
$\begin{array}[]{ @{} l @{} }x_{t}=c\left[y-\left(\displaystyle\frac{x^{3}}{3}-x\right)+z\right],\\ \\[3.0pt] y_{t}=-\displaystyle\frac{1}{c}(x-a+by).\end{array}$ (5)
FitzHugh introduced the parameter $z$ to represent membrane current density, the variable $x$ is related to the membrane voltage and the sodium activation, and the variable $y$ corresponds to the sodium inactivation and the potassium activation. In this context, these equations thus work as an activator-inhibitor model. Later, Nagumo, Arimoto, and Yoshizawa proved the equivalence of this model with a electrical circuit. Fig. 2A shows this electrical equivalent, which includes a capacitor (representing membrane capacitance $C$), a tunnel diode [$F(V)$], a resistor (representing channel resistance $R$), an inductor ($L$), and a battery ($E$). The dynamics of this electrical circuit are captured by the following equations:
$\begin{array}[]{ @{} l @{} }CV_{t}=I-F(V)-W,\\[3.0pt] LW_{t}=E-RW+V,\end{array}$ (6)
where $V$ is the voltage across the circuit, $W$ corresponds to the current through the $R-L-E$ branch of the circuit, and $F(V)$ is the current flowing through the tunnel diode driven by the voltage $V$. When the nonlinear function $F(V)$ is cubic, this set of equations can be readily transformed into Eqs. (5).
The model Eqs. (5) are invariant under the transformation $x\to-x$, $y\to-y$ and $a\to-a$, so considering $a>0$ is sufficient to understand its dynamics. The equations are also invariant under the transformation $(t,c)\to(-t,-c)$, so we can also restrict ourselves to $c>0$. To cast these equations in a simpler form, we consider the transformation $t\rightarrow ct$ and $\varepsilon=1/c^{2}$. We also absorb the parameter $z$ into the variable $y$ ($y\rightarrow y+z$), such that we redefine $a\rightarrow a-bz$:
$\begin{array}[]{ @{} l @{} }x_{t}=-\displaystyle\frac{x^{3}}{3}+x+y,\\ \\[3.0pt] y_{t}=\varepsilon(x+by+a).\end{array}$ (7)
@@
$\begin{array}[]{ @{} l @{} }u_{t}=-u^{3}+u-v,\\[3.0pt] v_{t}=\varepsilon(u-bv+a),\end{array}$ (8)
@@ I.3 Excitability in the generalized FitzHugh-Nagumo model
As observed in Fig. 2B, the main dynamics of the HH model are also observed in the FHN model, i.e. excitability and relaxation-like oscillations. Excitability is classified as type II, characterized by the lack of a distinct threshold. Excitability and oscillatory dynamics are essential to model neuronal spiking, and although the quantitative behavior may change, the qualitative behavior is preserved by the simplified model. The key ingredients that preserve these dynamical behaviors are the time scale separation and the cubic shape of the first differential equation. This separation implies the presence of fast and slow regions (quick versus slow changes in Fig. 2B), while the influence of the latter will be explained in detail in the next Section II, Subsection II.4.
@@
$\begin{array}[]{ @{} l @{} }u_{t}=f(u,v),\\[3.0pt] v_{t}=\varepsilon g(u,v),\end{array}$ (9)
where $f(u,v)$ is a nonlinear function of $u$ and a linear function of $v$, $g(u,v)$ a linear function of both $u$ and $v$, and $\varepsilon$ represents the time scale separation between both equations. The nonlinearity of $f(u,v)$ is typically represented by a cubic function, written as $-u^{3}+u$ as in Eq. (8) or factorized as $u(1-u)(u-c)$, which is equivalent because the quadratic term can be eliminated as proven in Rocsoreanu’s book Rocsoreanu et al. (2012). The cubic shape of $f(u,v)$ is also often approximated piecewise-linearly using $f(u,v)=-H(u-a)+u-v$ with $H$ a Heaviside function, which was first proposed by McKean in the 1970s McKean Jr (1970).
Figure 2: Electrical analog of the FitzHugh-Nagumo model A. Circuit representation: The FHN model’s electrical counterpart includes key components that mimic biological neuronal dynamics. This circuit comprises: (i) A capacitor, symbolizing the neuronal membrane’s capacitance $C$. (ii) A tunnel diode, representing the nonlinear ionic current $F(V)$. (iii) A resistor, indicative of the channel resistance $R$. (iv) An inductor $L$ and a battery $E$, completing the circuit to model the rest of the system’s dynamics. B. Neuronal dynamics regimes: Through simulations of Eq. (8), we capture essential neuronal behaviors: (i) Excitable regime: At $(a,b,\varepsilon)=(0.1,1.5,0.01)$, the circuit mimics the excitable nature of neurons, responding robustly to stimuli beyond a certain threshold. (ii) Oscillatory regime: With parameters $(a,b,\varepsilon)=(0,0.5,0.01)$, the system exhibits periodic oscillations, typical of active neuronal firing patterns. C. Spatial coupling in neuronal arrays: Extending the model to encompass spatial interactions involves linking multiple such circuits. This approach, as demonstrated in J. Nagumo et al.’s seminal work Nagumo et al. (1962), allows for the exploration of wave propagation and collective behaviors in a network of neuron-like elements, mirroring the complex dynamics observed in biological neural networks (extracted from Fig. 17 of the original paper Nagumo et al. (1962)).

### v4

@@ I Introduction
Figure 1: Exploring neuronal dynamics: The Hodgkin-Huxley model and its impact on subsequent research via the FitzHugh-Nagumo model. A. Depiction of the action potential across an axonal membrane, adapted from Hodgkin and Huxley’s seminal work Hodgkin and Huxley (1939), with the potential difference measured in millivolts and the external environment set as the zero potential reference. B. Illustration of ion channels within a neuronal axon, detailing the exchange of sodium and potassium ions between the extracellular medium (EM) and the intracellular medium (IM). C. Demonstration of excitable behavior as characterized by the HH model, adapted from Fig. 17 from the original study Hodgkin and Huxley (1952). D. Publication and citation data gathered from the Web of Science using “FitzHugh-Nagumo” as the search term (September 19, 2024). Note that these metrics do not include papers where the FHN model is not the primary focus, or papers published before the 1970s when the model name was first used. This shows the impact of the subsequent research using the FHN model as a simplified description of the HH model dynamics.
@@ I.1 Action potentials in neurons: the Hodgkin-Huxley model
Using voltage-clamp methods, they discovered that the permeability of the membrane for sodium and potassium was regulated independently, with the conductance depending on both time and the membrane potential. Remarkably, they provided a complete computational model for the action potential in a single cell, now known as the Hodgkin-Huxley (HH) model Hodgkin and Huxley (1952), which was a major advancement of our understanding. For their seminal work, they were awarded the Nobel Prize in Physiology or Medicine in 1963 Nob (1963).
@@
$\begin{array}[]{ @{} l @{} }I=C_{m}\frac{dV_{m}}{dt}+\overline{g}_{K}n^{4}(V_{m}-V_{k})+\overline{g}_{Na}m^{3}h(V_{m}-V_{Na})+\overline{g}_{l}(V_{m}-V_{l}),\vspace{3pt}\\ \displaystyle\frac{dn}{dt}=\alpha_{n}(V_{m})(1-n)-\beta_{n}(V_{m})n,\vspace{3pt}\\ \displaystyle\frac{dm}{dt}=\alpha_{m}(V_{m})(1-m)-\beta_{m}(V_{m})m,\vspace{3pt}\\ \displaystyle\frac{dh}{dt}=\alpha_{h}(V_{m})(1-h)-\beta_{h}(V_{m})h.\end{array}$ (1)
Here, $I$ represents the current per unit area, while $\alpha_{i}$ and $\beta_{i}$ denote voltage-dependent rate constants specific to the $i$th ion channel. $\overline{g}_{n}$ represents the maximum conductance value, and $n$, $m$, and $h$ are dimensionless probabilities ranging between 00 and $1$. These probabilities correspond to the activation states of potassium channel subunits, the activation of sodium channel subunits, and the inactivation of sodium channel subunits, respectively.
@@
The fact that the HH model correctly captures these excitable dynamics Hodgkin and Huxley (1952) was a major realization as it provided a detailed understanding of the measured action potential and the dynamical process of neural excitability (Fig. 1C).
@@ I.2 A simplified description of neuronal excitability: the FitzHugh-Nagumo model
FitzHugh started from the van der Pol oscillator equation Kanamaru (2007), introduced in the 1920s by Van der Pol Van (1927). As its name suggests, the Van der Pol oscillator admits oscillations which are relaxation-like, meaning they feature periods of slow progress close to a “low state” and a “high state” punctuated by rapid transitions between them. Van der Pol’s equation is built from the simple differential equation for the damped harmonic oscillator by replacing the damping constant with a damping coefficient that depends quadratically on $x$, thus introducing a nonlinearity:
@@
$\begin{array}[]{ @{} l @{} }x_{t}=c\left[y-\left(\displaystyle\frac{x^{3}}{3}-x\right)\right],\vspace{3pt}\\ y_{t}=-\displaystyle\frac{1}{c}x,\end{array}$ (4)
@@
$\begin{array}[]{ @{} l @{} }x_{t}=c\left[y-\left(\displaystyle\frac{x^{3}}{3}-x\right)+z\right],\vspace{3pt}\\ y_{t}=-\displaystyle\frac{1}{c}(x-a+by).\end{array}$ (5)
FitzHugh introduced the parameter $z$ to represent membrane current density, the variable $x$ is related to the membrane voltage and the sodium activation, and the variable $y$ corresponds to the sodium inactivation and the potassium activation. In this context, these equations thus work as an activator-inhibitor model. Later, Nagumo, Arimoto, and Yoshizawa proved the equivalence of this model with an electrical circuit. Fig. 2A shows this electrical equivalent, which includes a capacitor (representing membrane capacitance $C$), a tunnel diode [$F(V)$], a resistor (representing channel resistance $R$), an inductor ($L$), and a battery ($E$). The dynamics of this electrical circuit are captured by the following equations:
$\begin{array}[]{ @{} l @{} }CV_{t}=I-F(V)-W,\vspace{3pt}\\ LW_{t}=E-RW+V,\end{array}$ (6)
where $V$ is the voltage across the circuit, $W$ corresponds to the current through the $R-L-E$ branch of the circuit, and $F(V)$ is the current flowing through the tunnel diode driven by the voltage $V$. When the nonlinear function $F(V)$ is cubic, this set of equations can be readily transformed into Eqs. 5.
Figure 2: Electrical analog of the FitzHugh-Nagumo model A. Circuit representation: The FHN model’s electrical counterpart includes key components that mimic biological neuronal dynamics. This circuit comprises: (i) A capacitor, symbolizing the neuronal membrane’s capacitance $C$. (ii) A tunnel diode, representing the nonlinear ionic current $F(V)$. (iii) A resistor, indicative of the channel resistance $R$. (iv) An inductor $L$ and a battery $E$, completing the circuit to model the rest of the system’s dynamics. B. Neuronal dynamics regimes: Through simulations of Eq. 8, we capture essential neuronal behaviors: (i) Excitable regime: At $(a,b,\varepsilon)=(0.1,1.5,0.1)$, the circuit mimics the excitable nature of neurons, responding robustly to stimuli beyond a certain threshold. (ii) Oscillatory regime: With parameters $(a,b,\varepsilon)=(0,0.5,0.01)$, the system exhibits periodic oscillations, typical of active neuronal firing patterns. C. Spatial coupling in neuronal arrays: Extending the model to encompass spatial interactions involves linking multiple such circuits. This approach, as demonstrated in J. Nagumo et al.’s seminal work Nagumo et al. (1962), allows for the exploration of wave propagation and collective behaviors in a network of neuron-like elements, mirroring the complex dynamics observed in biological neural networks.
The model Eqs. 5 are invariant under the transformation $x\to-x$, $y\to-y$ and $a\to-a$, so considering $a>0$ is sufficient to understand its dynamics. The equations are also invariant under the transformation $(t,c)\to(-t,-c)$, so we can also restrict ourselves to $c>0$. To cast these equations in a simpler form, we consider the transformation $t\rightarrow ct$ and $\varepsilon=1/c^{2}$. We also absorb the parameter $z$ into the variable $y$ ($y\rightarrow y+z$), such that we redefine $a\rightarrow a-bz$:
$\begin{array}[]{ @{} l @{} }x_{t}=-\displaystyle\frac{x^{3}}{3}+x+y,\vspace{3pt}\\ y_{t}=\varepsilon(x+by+a).\end{array}$ (7)
@@
$\begin{array}[]{ @{} l @{} }u_{t}=-u^{3}+u-v,\vspace{3pt}\\ v_{t}=\varepsilon(u-bv+a),\end{array}$ (8)
@@ I.3 Excitability in the generalized FitzHugh-Nagumo model
As observed in Fig. 2B, the main dynamics of the HH model are also observed in the FHN model, i.e. excitability and relaxation-like oscillations. Excitability is classified as type II, characterized by the lack of a distinct threshold. Excitability and oscillatory dynamics are essential to model neuronal spiking, and although the quantitative behavior may change, the qualitative behavior is preserved by the simplified model. The key ingredients that preserve these dynamical behaviors are the time scale separation and the cubic shape of the first differential equation. This separation implies the presence of fast and slow regions (quick versus slow changes in Fig. 2B), while the influence of the latter will be explained in more detail in Fig. II.4.
@@
$\begin{array}[]{ @{} l @{} }u_{t}=f(u,v),\vspace{3pt}\\ v_{t}=\varepsilon g(u,v),\end{array}$ (9)
where $f(u,v)$ is a nonlinear function of $u$ and a linear function of $v$, $g(u,v)$ a linear function of both $u$ and $v$, and $\varepsilon$ represents the time scale separation between both equations. The nonlinearity of $f(u,v)$ is typically represented by a cubic function, written as $-u^{3}+u$ as in Eq. 8 or factorized as $u(1-u)(u-c)$, which is equivalent because the quadratic term can be eliminated as shown in Rocsoreanu’s book Rocsoreanu et al. (2012). The cubic shape of $f(u,v)$ is also often approximated piecewise-linearly using $f(u,v)=-H(u-a)+u-v$ with $H$ a Heaviside function, which was first proposed by McKean in the 1970s McKean (1970).